Solve the following differential equations cosx d²y/dx²+sinx dy/dx - 2ycos³x = 2cos^5 x
1
Expert's answer
2020-07-16T19:25:21-0400

"\cos xy''+\sin xy'-2y\cos^3x=2\cos^5x"

Using a substitution

"y=-\cos^2x+1+z(x)\\
y'=\sin2x+z'(x)\\
y''=2\cos2x+z''(x)"

Then

"\cos x(2\cos2x+z'')+\sin x(\sin2x+z')-\\
-2(-\cos^2x+1+z)\cos^3x=2\cos^5x\\
\cos xz''+\sin xz'-2\cos^3xz=0"

Using a substitution

"\sin x=t\\
x=\sin^{-1}t\\
z'=\frac{z'_t}{x'_t}=z'\sqrt{1-t^2}\\
z''=\frac{z''_t}{x'_t}=(z''(1-t^2)+z'\frac{-2t}{\sqrt{1-t^2}})\sqrt{1-t^2}\\"

Put it into the equation

"(1-t^2)^{\frac{3}{2}}(z''-2z)=0\\
z''-2z=0\\
\lambda^2-2=0\\
\lambda_1=\sqrt2\\
\lambda_2=-\sqrt2\\
z=C_1e^{\sqrt2 t}+C_2e^{-\sqrt2 t}\\
z=C_1e^{\sqrt2 \sin x}+C_2e^{-\sqrt2 \sin x}\\
y=-\cos^2 x+1+C_1e^{\sqrt2 \sin x}+C_2e^{-\sqrt2 \sin x}\\"